A car travels at 50 miles per hour and can go 300 miles between refueling stops. If each refueling takes an hour and the car starts with a full tank, how long will a 1000-mile journey take?
Answer by addingup(3677)

23 hours, like this:
300/50 = 6 hours between refills+1 hour to refill = 7 hours. Means:
- refill #1 at 300 miles
- refill #2 at 600 miles
- refill #3 at 900 miles
3 refills times 7 = 21 hours. Plus:
1000-900 = 100
100/50mph (the car's speed) = 2 additional hours
21+2 = 23 hours total for this journey.